OpenStudy (anonymous):

Im doing some physics and i cant figure out this problem... A karate expert breaks a stack of bricks with his bare hand. If the force applied is 520 newtons and the impact time is 0.5 × 10ȳ seconds, what is the value of impulse?

OpenStudy (anonymous):

\[J=\Delta p\]=\[p2-p1=F\Delta t=520*50s =\] that is if your time is \[.5*10^3\] if its \[.5*10^{-3}\] \[.0005

OpenStudy (anonymous):

i think it's supposed to be .0005 because 50 seconds is a huge impact time, it's usually less than a second

OpenStudy (anonymous):

this is why they show the impacts of sports and other stuff in slow motion... it's way to hard to determine the impact timing